OpenStudy (anonymous):

If y varies directly as x, and y is 18 when x is 5, which expression can be used to find the value of y when x is 11?

OpenStudy (jhannybean):

direct variation: y = kx given y=18, x=5, find k. k = y/x k = (18)/5 = ?

OpenStudy (anonymous):

what would be the answer?

OpenStudy (jhannybean):

you'll find out once you solve the problem :P

OpenStudy (jhannybean):

what does k equal?

OpenStudy (anonymous):

3.6

OpenStudy (jhannybean):

Now you use k=3.6 , x=11 into the formula once again. y = kx = (3.6)(11) =?

OpenStudy (anonymous):

39.6

OpenStudy (jhannybean):

There you go :)

OpenStudy (anonymous):

my answer choices are 1/9, 2/3/,54,96 though :(

OpenStudy (jhannybean):

Something is wrong with this question, or your answers.